OpenStudy (blutz):

two side of a traingular plot of groud are 256.8 feet and 195.2 feet respectively the included angle being 65 mins 27 seconds. Find(a) the length of fence required to enclose the plot,(b) the area of the plot

KjSaif (kjsaif):

should there be a picture?

KjSaif (kjsaif):

you there @blutz?

OpenStudy (owen3):



OpenStudy (owen3):

1 degree = 60 minutes 1 minute = 60 seconds

OpenStudy (owen3):

First we have to convert minutes and seconds to degrees. 65 mins 27 seconds= 65/60 + 27/60^2 degrees = 1.090833333 degrees Using law of cosines I get that the third side is 61.747 feet The total fencing you need is the perimeter of the triangle 61.747 +256.8 + 195.2= 513.747 feet The area is 1/2 * 195.2 * 256.8 * sin ( 1.090833333) = 477.15 ft^2

OpenStudy (owen3):



OpenStudy (owen3):

That picture is more accurate.

OpenStudy (jango_in_dtown):

seems correct.